Fermi National Laboratory


Accelerator Update

Monday March 22
The day shift began with Antiproton Source (Pbar) experts conducting lattice studies, with operations preparing to send beam to MiniBooNE, and with the Tevatron (TeV) work continuing.

At 9 AM, Pbar experts switched their studies to reverse protons.

Operations established beam to MiniBooNE for multiwire testing at 9:11 AM. Operators turned off MiniBooNE beam at 10:25 AM to prepare the area for an access.

Operations turned off the Pbar beam for an access at 12:56 PM.

A Linac expert finished tuning the beam line at 1:49 PM.

Operators completed interlocking the D-zero (D0) collision hall at 2:38 PM.

At 3:32 PM, a Neutron Therapy Facility (NTF) study began causing problems with the Linac beam. Linac and NTF experts tuned a correction magnet and had the problem resolved within fifteen minutes.

Vacuum technicians accessed the F1 Switchyard (SY) enclosure to search for a vacuum leak at 3:36 PM.

A Pbar crate stopped communicating earlier this afternoon. It was noticed at 3:50 PM because of a Debuncher cryo system vacuum problem. Operators called the Cryo Coordinator for advice, rebooted the crate, and began recovering the cryo and vacuum systems.

Vacuum techs reported at 6:16 PM that they had found the SY vacuum leak and worked on the problem for about one and a half hours.

By 7:39 PM, operators had interlocked TeV sector B and the Transfer Hall.

A Recycler expert began a three-hour study period at 7:45 PM.

A SY expert began a study period at 9 PM.

Operators completed interlocking Meson enclosures M02 and M03 at 9:54 PM.

Tuesday March 23
The midnight shift began with Pbar conducting reverse proton studies, and with SY experts conducting dump studies.

At 1:49 AM, SY experts completed their studies.

Recycler studies began at 7:35 AM.

Cryo experts began warming up the TeV sector B4 cryo system at 10:15 AM because of contamination concerns.

A Linac intertank phase problem for drift tube tank #2 reappeared at 10:47 AM. Then, at 11:15 AM, drift tube tank #1 caused the Linac to trip for the first of four times in ten minutes.

At 11:34 AM, a Preaccelerator expert did some preliminary work, which held off Linac beam, in order to take down the I- Source in the near future.

Linac experts began changing a module for a Linac RF station (LRF1) at 11:38 AM.

Linac beam resumed at 12:07 PM.

Pbar began stacking at 12:27 PM.

Operations received permission at 4:47 PM to turn on the TeV after performing a successful hipot.

At 6:16 PM, operators completed interlocking the D-zero (D0) collision hall.

NTF began taking Linac beam at 6:47 PM.

A TeV expert reported at 7:22 PM that all work at TeV sector A4 had been completed and the cryo system cooldown could begin.

Two Booster RF stations (BRF7 & 8) tripped off on high vacuum pressure at 7:44 PM. They both reset okay.

SY studies began at 9 PM.

Operators called in two power supply techs to work on an AP1 transfer line power supply at 9:25 PM. They had the station back up by 10:12 PM.

A problem developed with MI/SY magnet (HV100) at 10:24 PM.

Both NTF and Recycler studies had ended by 10:55 PM.

At 11:24 PM, a TeV expert said the TeV sector B1 cryo system could begin cooldown.

Wednesday March 24
The midnight shift began with Operations preparing to access the TeV sector F0 to investigate HV100.

At 12:45 AM, techs and experts accessed TeV sector F0. A tech had found a problem (no load) with HV100 before they began the access. At 1:08 AM, they found a damaged connector and a LCW leak. Repairs began immediately.

Operators completed interlocking TeV sectors A - E at 2:35 AM.

Operators completed interlocking the Collider Detector at Fermilab (CDF) collision hall at 3:43 AM.

Work on MV100 finished and all techs were out of F sector by 5:30 PM.

Operators completed interlocking the TeV sector C0 collision hall at 5:53 AM.

Operators successfully hipotted the TeV at 5:57 AM.

Operators reported at 6:17 AM that they were having problems racking in feeder 23 at the Master Substation. The Duty Electrician investigated and then called in the high voltage experts.

Pbar began stacking at 7:48 AM.

Today's Plans
The main plan for today is to complete the TeV alignment work in the CDF collision hall. During this work, Pbar will stack and conduct studies along with the other machines. After the CDF alignment work finishes sometime later this evening, TeV experts will begin TeV commissioning and then go into a study period.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/24/2004   email Fermilab